OpenStudy (anonymous):

Mai borrowed $12,000 at 5% interest for 6 years. What was the total interest Mai paid? $

OpenStudy (anonymous):

so can u help me plz

OpenStudy (anonymous):

k

OpenStudy (anonymous):

it is not nearly that simple are you given any formulas to compute this with?

OpenStudy (anonymous):

3600

OpenStudy (anonymous):

yeah there's a lil more to it than that

OpenStudy (anonymous):

actually there is a ton more to it

OpenStudy (anonymous):

no not really because all you need is the total amount of interest he paid not the total amount

OpenStudy (anonymous):

actually doing what mehek said will get you the answer to your question

OpenStudy (anonymous):

like hell

OpenStudy (anonymous):

prove me wrong cuz you can't

OpenStudy (anonymous):

you need to know your monthly payments then you need to calculate the total amount of money paid then you can figure out what the total interest is

OpenStudy (anonymous):

CRAP Sorry focusing on multiple things right now.

OpenStudy (anonymous):

but the problem is yo don't know when he pays his bill

OpenStudy (wolf1728):

$12,000 at 5% interest for 6 years Here is the loan payment formula: http://www.1728.org/loanform.htm Monthly payment = rate + [rate]/([1+rate]^months]-1) * principal rate = 5/1200 = 0.004166667 Monthly payment = 0.004166667 + [0.004166667]/[(1.004166667)^72)-1] * principal Monthly payment = 0.004166667 + [0.004166667 / (1.3490177764 -1)] * 12,000 Monthly payment = 0.004166667 + [0.004166667 / .3490177764] * 12,000 Monthly payment = (0.004166667 + 0.0119382658) * 12,000 Monthly payment = 0.0161049328 * 12,000 Monthly payment = 193.2591941643 Monthly payment = 193.26 Total Payments = 193.26 * 72 months = $13,914.72 Total Interest Paid = $13,914.72 minus $12,000 $1,914.72 As satellite73 said, it is NOT that simple. (Here is a link to a calculator that will compute the monthly payment for you.) http://www.1728.org/calcloan.htm